# Quantization Meets Reasoning: Exploring LLM Low-Bit Quantization Degradation for Mathematical Reasoning

### v1

@@ Abstract
Large language models have achieved significant advancements in complex mathematical reasoning benchmarks, such as MATH. However, their substantial computational requirements present challenges for practical deployment. Model quantization has emerged as an effective strategy to reduce memory usage and computational costs by employing lower precision and bit-width representations. In this study, we systematically evaluate the impact of quantization on mathematical reasoning tasks. We introduce a multidimensional evaluation framework that qualitatively assesses specific capability dimensions and conduct quantitative analyses on the step-by-step outputs of various quantization methods. Our results demonstrate that quantization differentially affects numerical computation and reasoning planning abilities, identifying key areas where quantized models experience performance degradation.
@@
Zhen Li1thanks: These authors contributed equally to this work., Yupeng Su211footnotemark: 1, Runming Yang3, Zhongwei Xie4, Ngai Wong5thanks: Corresponding authors.hongxia.yang@polyu.edu.hk, Hongxia Yang122footnotemark: 2, 1The Hong Kong Polytechnic University, 2Southern University of Science and Technology, 3Tsinghua University, 4Wuhan University, 5The University of Hong Kong
1 Introduction and Related Work
@@
Figure 1: An overview of our proposed approaches. This example demonstrates the prediction differences before and after quantization, and that after format alignment and knowledge infilling through supervised fine-tuning. Finally, we analyze the impact at each step and categorize the types of error.
However, such advancements come at a cost. The computational requirements of LLMs—both in terms of memory and latency—pose significant practical barriers  Kaplan et al. (2020); Hoffmann et al. (2022); Gou et al. (2024). To address these efficiency challenges, researchers have explored model compression techniques such as pruning LeCun et al. (1989); Han et al. (2015), knowledge distillation  Hinton (2015); Jiao et al. (2019); Yang et al. (2024), and more recently, quantization  Hubara et al. (2018); Jacob et al. (2018); Yao et al. (2022). Quantization reduces memory usage and computational overhead by representing weights and activations in low-bit formats (e.g., INT8), halving GPU memory consumption and nearly doubling throughput in operations like matrix multiplication and attention  Rastegari et al. (2016); Lin et al. (2015); Dettmers et al. (2022). While it performs well on standard NLP tasks with minimal performance loss  Ma et al. (2024), its effect on complex mathematical reasoning—requiring precise, contextually coherent, and logical steps—remains unclear, particularly for tasks like MATH or Code.
@@
Against this backdrop, the interplay between quantization, advanced inference strategies, and the underlying reasoning fidelity of LLMs emerges as a key research question. To summarize, in this study, we present the following key contributions:
  • We evaluate quantization techniques across diverse models and tasks. Our findings show significant accuracy degradation on complex benchmarks like MATH, where all methods result in notable performance losses.
  • We introduce a method aligning step-by-step outputs during training, ensuring consistent knowledge exposure across full-precision and quantized models. This enables detailed reasoning while preserving format alignment.
  • To overcome limitations of benchmarks focused only on final answer correctness, we propose a framework analyzing mathematical reasoning across seven dimensions. Through qualitative and quantitative analyses, we identify specific areas where quantized models struggle in complex reasoning tasks.
@@ 2.1 Model Quantization
We investigate the effects of both weight-only quantization like GPTQ Frantar et al. (2022), AWQ Lin et al. (2024) and weight-activation quantization like SmoothQuant Xiao et al. (2023). We evaluate these methods under different quantization settings, including a weight precision of 4 bits combined with 16-bit activations (W4A16) and both weight and activation precision are 8 bits (W8A8). By employing mainstream quantization techniques, we provide a fair analysis of the quantized models.
2.2 Format Alignment & Knowledge Infilling
We employ a stage of format alignment and knowledge infilling to ensure the model internalizes the introduced tokens and responds effectively to instruction-following prompts. Specifically, we employ LoRA Hu et al. (2021); Wu et al. (2024) as a supervised fine-tuning (SFT) technique for quantized models. LoRA injections enable efficient refinement of a small subset of parameters, facilitating fine-tuning without extensive modifications to the entire parameter space.
We jointly train the embedding and output layers alongside the LoRA adapters to integrate new tokens and align them with the model’s underlying representation space. Since these tokens explicitly define reasoning steps and answer boundaries, ensuring their accurate interpretation is crucial. By combining LoRA adaptation with standard fine-tuning, we effectively “inject” knowledge of these stepwise structures into the model’s latent space. This not only enhances the model’s ability to process structured instructions but also improves its robustness against the challenges posed by quantization. Additionally, this approach allows us to better analyze the differences in inference steps before and after model quantization.
2.3 Data
We use PRM800K dataset Lightman et al. (2023), a large-scale resource containing step-level correctness annotations of model-generated solutions to problems from the MATH benchmark Hendrycks et al. (2021). These annotations allow models to learn granular reasoning steps, rather than simply producing final answers. To make the problem tractable within our limited computation and space, we sample 8,000 problem-solution pairs from PRM800K. Each sample provides a well-documented reasoning trajectory, enabling the model to learn how to follow instructions and accurately traverse multi-step logical derivations.
To facilitate stepwise reasoning Birchlabs , we introduce four special tokens: <|step_start|>, <|step_end|>, <|answer_start|>, and <|answer_end|>. These tokens serve as structured markers that guide the model through the reasoning process, from the initiation of a reasoning step to its conclusion, and from the start of an answer to its completion. By adding this explicit structure, we provide the model with a clearer signal for identifying the boundaries of reasoning steps and final answers. This structured annotation is crucial, as it sets the stage for the model to reliably follow instructions, even when compressed or quantized.
3 Experiments
We evaluate our models on the MATH dataset using OpenCompass evaluation framework Contributors (2023). Notably, for the Fine-tuned model, we use zero shot evaluation with prompts consistency between the prompts used during evaluation and those introduced during the knowledge injection phase. In contrast, vanilla pretrained models and quantized models are evaluated with few shot prompts to compensate for their limited instruction-following capability, aligning with the characteristics of pretrained models. This setup ensures a fair comparison across all models.
Model Llama-3.2-3B Llama-3.1-8B
Bit
@@ 3 Experiments
SQ
@@
SQ
Base SFT Base SFT Base SFT Base Base SFT Base SFT Base SFT Base
Score 5.62 9.04 3.88 6.46 4.64 5.22 5.28 7.70 15.30 5.30 11.56 4.98 8.64 6.86
$\Delta$downarrow1.74 $\downarrow$2.58 $\downarrow$0.98 $\downarrow$3.82 $\downarrow$0.34 - - $\downarrow$2.4 $\downarrow$3.74 $\downarrow$2.72 $\downarrow$6.66 $\downarrow$0.84
Table 1: This table presents the quantitative results on the MATH benchmark for vanilla BFloat16 and quantized models using Base and SFT methods. The $\Delta$ values indicate the performance degradation caused by quantization. Notably, the quantized models exhibit varying degrees of accuracy loss both before and after fine-tuning, indicating that quantization leads to a decline in the models’ ability to perform complex mathematical reasoning tasks.
3.1 General Setup
All experiments in this study were conducted on a cluster of 8 NVIDIA 80G A800 GPUs with a global batch size of 16. We applied LoRA fine-tuning to the vanilla model, while QLoRA Dettmers et al. (2024) was employed for the quantized models. Training was also performed on the embed_tokens and lm_head layers, using a learning rate of 1.0e-5. A cosine learning rate schedule with a warmup ratio of 0.05 was adopted for optimization.
3.2 Prompting
Figure 2 illustrates an example of the prompt utilized during the format alignment and knowledge infilling phase and evaluation phase. The prompt structure was carefully designed to align with the training inputs, ensuring that the models received consistent and clear instructions.
3.3 Results
@@
Table 1 compares the performance of full precision and quantized models on the MATH dataset for Llama-3.2-3B and Llama-3.1-8B. Both pretrain and finetune methods are evaluated. The comparison includes both the pretrain and finetune methods, measuring the performance (Score) and the corresponding loss ($\Delta$) introduced by quantization. The Vanilla model serves as the upper baseline, while quantized models exhibit varying degrees of accuracy loss. Finetune consistently improves performance across all settings. Among the quantized methods, SmoothQuant (W8A8) shows the smallest degradation, demonstrating better robustness compared to AWQ and GPTQ.
3.4 Case Study and Analyses
To investigate the underlying reasons for the performance degradation observed in the quantized models, we categorized the errors into seven distinct types:
  • Misunderstanding of Problem: Fails to accurately interpret the problem statement, leading to incorrect assumptions or interpretations.
  • Logical Error: Flaws in the reasoning process, where the sequence of steps lacks logical coherence or correctness.
  • Computation Error: Mistakes in arithmetic or algebraic calculations within the steps.
  • Formula Misuse: Incorrect application or selection of mathematical formulas and theorems relevant to solving the problem.
  • Step Omission: Missing crucial steps in the reasoning process, resulting in incomplete or incorrect solutions.
  • Boundary Condition Error: Inadequate consideration of special cases or boundary conditions that affect the problem’s outcome.
  • Symbol Error: Incorrect usage of mathematical symbols, such as misplacing signs or using incorrect variables.
Figure 3: The relative performance loss of different quantization methods across capability dimensions is measured with the full-precision model’s performance normalized to a baseline value of 1.0.
Using the defined error_types_list, we leveraged the GPT-4o API to perform a quantitative analysis of the quantized models’ errors. The process involved the following steps:
  • Automated Error Classification: For each erroneous output from the quantized models, the GPT-4o API was prompted to classify the error based on the predefined error types.
  • Manual Verification: To ensure the accuracy of the automated classifications, a subset of the results was manually reviewed and helped in refining the prompt for better precision.
@@ 3.4 Case Study and Analyses
The Figure 3 & 4 reveal that computation errors and step omissions are the most significant failure modes in quantized models. Computation errors, particularly prominent in AWQ-W4A16, are likely caused by low-bit precision overflow and underflow, which propagate inaccuracies across multi-step calculations. SmoothQuant-W8A8 shows a notable increase in step omissions, likely due to reduced activation precision, which disrupts the retention of intermediate reasoning steps. While boundary condition errors are less frequent, this may reflect a “survivor effect”, where earlier errors, such as logical or computation failures, prevent the models from reaching steps where boundary conditions would be tested. These results suggest that quantization disrupts step-by-step reasoning, exposing the models’ limitations in handling complex mathematical tasks. The underlying causes of these errors may stem from the limited numerical precision and accumulated quantization noise in low-bit formats, which impair the representation of long-term dependencies and intermediate values in complex reasoning tasks.
Figure 4: Comparative analysis of error types for quantized models (AWQ-W4A16, GPTQ-W4A16, and SmoothQuant-W8A8).

### v2

@@ Abstract
Large language models have achieved significant advancements in complex mathematical reasoning benchmarks, such as MATH. However, their substantial computational requirements present challenges for practical deployment. Model quantization has emerged as an effective strategy to reduce memory usage and computational costs by employing lower precision and bit-width representations. In this study, we systematically evaluate the impact of quantization on mathematical reasoning tasks. Our results demonstrate that aggressive quantization methods like AWQ and GPTQ introduce up to 32.39% accuracy degradation (average 11.31%) on Llama-3 models, particularly in numerical computation and reasoning planning. To address this, we introduce a multidimensional evaluation framework combining qualitative capability analysis and quantitative error assessment. We further develop targeted recovery strategies, showing that fine-tuning quantized models on only 545 task-specific examples for 3 minutes on 4 GPUs effectively restores reasoning capabilities to near full-precision levels. Additionally, our error assessment pipeline achieves 98.9% accuracy in diagnosing and localizing errors across 3,366 failure cases, providing actionable insights for mitigating quantization-induced degradation.
@@
Zhen Li1thanks: These authors contributed equally to this work., Yupeng Su211footnotemark: 1, Runming Yang3, Congkai Xie4, Zheng Wang3, Zhongwei Xie5, Ngai Wong6thanks: Corresponding authors: Ngai Wong(nwong@eee.hku.hk), Hongxia Yang(hongxia.yang@polyu.edu.hk), Hongxia Yang122footnotemark: 2, 1Hong Kong Polytechnic University, 2Southern University of Science and Technology, 3Tsinghua University, 4Reallm Labs, 5Wuhan University, 6University of Hong Kong
1 Introduction
@@
However, such advancements come at a cost. The computational requirements of LLMs, both in terms of memory and latency, pose significant practical barriers  Kaplan et al. (2020); Hoffmann et al. (2022); Gou et al. (2024). To address these efficiency challenges, researchers have explored model compression techniques such as pruning LeCun et al. (1989); Han et al. (2015), knowledge distillation  Hinton (2015); Jiao et al. (2019); Yang et al. (2024), and more recently, quantization  Hubara et al. (2018); Jacob et al. (2018); Yao et al. (2022). Quantization reduces memory usage and computational overhead by representing weights and activations in low-bit formats (e.g., INT8), halving GPU memory consumption and nearly doubling throughput in operations like matrix multiplication and attention  Rastegari et al. (2016); Lin et al. (2015); Dettmers et al. (2022). While it performs well on standard NLP tasks with minimal performance loss  Ma et al. (2024), its effect on complex mathematical reasoning, requiring precise, contextually coherent, and logical steps, remains unclear, particularly for tasks like MATH or Code.
Figure 1: Pipeline of our study for evaluating and restoring reasoning capabilities in quantized models. By format alignment training and our error assessment pipeline with expert judge models, we diagnose and analyze ability-level reasoning weaknesses on model’s step-by-step solutions. Based on identified error types, we sample a targeted ‘medicine’ dataset to fine-tune the model via Direct Preference Optimization (DPO), aiming to restore performance while preserving efficiency.
@@
Against this backdrop, the interplay between quantization, advanced inference strategies, and the underlying reasoning fidelity of LLMs emerges as a key research question. In this work, we seek to illuminate these relationships, providing insights into how quantization influences mathematical reasoning and how we might mitigate these effects.
To summarize, in this study, we present the following key contributions:
  • We reveal that mainstream quantization methods (AWQ, GPTQ) incur substantial reasoning degradation, with up to 32.39% accuracy loss (average 11.31%) on MATH for Llama-3 models, exposing critical vulnerabilities in low-bit numerical representations.
  • We propose a step-aligned evaluation protocol to dissect reasoning errors into four dimensions (conceptual, methodological, executional, logical), enabling granular diagnosis of quantization effects. Our automated error assessment pipeline achieves 98.9% accuracy in categorizing 3,366 failure cases, surpassing human-in-the-loop baselines.
  • We demonstrate that quantized models recover the performance gaps through lightweight fine-tuning on only 545 targeted examples, requiring only 3 minutes of training, thus enabling efficient deployment without sacrificing reasoning fidelity.
2 Related Work
2.1 Quantization Techniques
Modern quantization methods balance efficiency and performance through minimizing the model output differences after quantization. Post-training quantization (PTQ) approaches Frantar et al. (2022); Lin et al. (2024); Xiao et al. (2023); Yao et al. (2022) enable efficient compression without retraining, while quantization-aware training (QAT) methods Hu et al. (2021); Dettmers et al. (2024) preserve task-specific performance through learnable scaling factors. Recent work extends these techniques to LLMs, though primarily evaluated on language understanding rather than reasoning tasks.
2.2 Mathematical Reasoning in LLMs
The emergence of specialized models like Minerva Lewkowycz et al. (2022) demonstrates LLMs’ potential for advanced mathematical problem-solving. Chain-of-Thought prompting Wei et al. (2022) and its variants (e.g., Program-of-Thought Chowdhery et al. (2023)) enhance multi-step reasoning by decomposing problems into interpretable sub-steps. However, these approaches assume high-precision model representations, potentially conflicting with quantization’s reduced numerical precision.
2.3 Model Alignment and Reasoning Optimization
Recent advances in model alignment integrate instruction tuning, pioneered by the FLAN framework Wei et al. (2021) and popularized through Alpaca Taori et al. (2023), with preference optimization techniques like DPO Rafailov et al. (2024). Concurrently, reasoning reliability enhancements employ self-consistency voting Wang et al. (2022) and process-based reward models Lightman et al. (2023a), building upon foundational work in verifiable reasoning Creswell and Shanahan (2022). Our investigation extends these directions by analyzing how quantization affects: the instruction-following capabilities crucial for step-by-step reasoning, and the consensus-building mechanisms in ensemble-based reasoning methods.
@@ 3.1 Model Quantization
We conduct a comprehensive investigation into the effects of quantization techniques, examining both weight-only quantization methods (GPTQ Frantar et al. (2022), AWQ Lin et al. (2024)). Our evaluation encompasses various quantization configurations, specifically focusing on 4-bit weight precision with 16-bit activations (W4A16). Through the systematic application of these mainstream quantization techniques, we provide a rigorous and balanced analysis of the resulting quantized models, offering valuable insights into their performance characteristics and trade-offs.
3.2 Format Alignment
To address the challenge of inconsistent instruction following and irregular output formatting in model-generated solutions, we introduce a format alignment stage. This phase aims to instill in the model a structured, step-by-step reasoning workflow without altering its underlying mathematical knowledge. Crucially, the objective here is not to teach the model new mathematical facts, but rather to ensure strict adherence to a prescribed output format, thereby enabling reliable qualitative and quantitative analysis of reasoning capability across quantized and full-precision variants.
We employ LoRA Hu et al. (2021) and QLoRA Dettmers et al. (2024) for full-precision model and quantized model respectively as lightweight adaptation techniques for format alignment. By selectively updating a small subset of parameters, these methods efficiently align knowledge of step-by-step solution formats into the model’s latent space without extensive retraining. This fine-tuning enables us to observe how multi-step reasoning is preserved or altered once the model is quantized, offering deeper insights into any capability loss induced by compression.
MATH
@@ 3.2 Format Alignment
Llama-3.1-8B-Instruct score 47.2 41.8 43.06
$\Delta$downarrow5.4(11.44%) $\downarrow$5.8(12.29%)
Llama-3.2-3B-Instruct score 40.2 38.8 31.6
$\Delta$downarrow2.16(4.97%) $\downarrow$4.0(10.42%)
Llama-3.2-1B-Instruct score 18.6 14.26 14.94
$\Delta$downarrow2.8(16.67%) $\downarrow$6.8(32.39%)
Table 1: Evaluation results of full-precision models and 4-bit weight, 16-bit activation quantization methods (AWQ and GPTQ) on the MATH benchmark. All three model scales exhibit varying degrees of performance degradation, posing challenges in solving mathematical reasoning problems. Notably, smaller-scale models suffer a performance drop exceeding 30%.
For training, we utilize the PRM800K dataset Lightman et al. (2023b), which provides 800K step-level correctness annotations from 75K solutions to 12K problems. These annotations supply granular, step-by-step reasoning trajectories, equipping models to separate complex problem-solving processes into well-defined stages. To reinforce this structure, we adopt a consistent system prompt across training and evaluation, ensuring that the boundaries of logical steps and final answers are clearly delineated. This consistent, step-by-step alignment is a necessary foundation for our subsequent qualitative and quantitive analyses of quantization-induced degradation in mathematical reasoning.
3.3 Analyses for Reasoning Errors
To investigate the underlying reasons for the performance degradation observed in the quantized models, we categorized the errors into seven specific error types in four major error categories.:
  • Conceptual Errors
    • Conceptual misunderstanding: Inadequate grasp of fundamental concepts or principles, leading to an incorrect framing of the problem or misguided approach to the solution.
    • Contextual oversight: Failure to account for relevant contextual factors or domain constraints (e.g., physical limits, geometric considerations) that significantly affect the problem’s outcome.
  • Method Errors
    • Procedural error: Errors in executing or adhering to prescribed procedures (e.g., failure to follow a standard algorithm or overlooking a required step), resulting in incomplete or invalid solutions.
    • Formula rule error: Improper application of mathematical rules, theorems, or formulae (including misapplication of a formula to an inappropriate scenario), which undermines the validity of the derived results.
  • Execution Errors
    • Computational error: Mistakes in arithmetic or algebraic manipulation, such as incorrect summation, factorization, or expansion, that compromise the accuracy of the final answer.
    • Symbolic manipulation error: Inaccurate handling of symbols or expressions, including mislabeling variables or misinterpreting symbolic transformations, leading to an incorrect representation of the problem.
  • Reasoning Errors
    • Logical reasoning error: Breakdown in the logical chain of thought, where the inference steps do not follow coherently or omit essential logical links, causing a disconnect between premises and conclusion.
3.4 Capability Restore
Building upon our systematic error analysis, we develop targeted capability restoration strategies to address quantization-induced performance degradation. For each identified error category (conceptual, methodological, executional, and logical), we create focused training subsets by pairing failed step-wise responses with correct responses to form a dataset of positive and negative pairs:
  • Conceptual Restoration: High-context problems requiring domain-specific constraints.
  • Methodological Restoration: Multi-step derivations with annotated procedure adherence.
  • Executional Restoration: Computation-intensive problems with stepwise verification.
  • Logical Restoration: Chain-of-thought reasoning with explicit inference links.
Figure 3: Distribution of error types in quantized models, highlights the dominant error types affecting mathematical reasoning in quantized models.
Figure 4: The radar plot of error distributions across different scale Llama models on the MATH benchmark.
We implement Direct Preference Optimization (DPO) with parameter-efficient adaptation, using QLoRA for quantized models. This dual approach enables simultaneous optimization of two objectives: (1) maximizing the likelihood of correct solution trajectories through supervised fine-tuning on error-specific subsets, and (2) addressing the model’s capability gaps while leveraging the undiminished foundational abilities of the quantized model to complete the solution, thereby enhancing overall mathematical reasoning performance.
Our recovery process comprises three phases:
  1. 1.
    Data Extraction and Construction: We extract and construct a dataset from the quantized models’ failed responses across different model scales.
  2. 2.
    Data Deduplication and Cross-Scale Sampling: To avoid redundancy, we deduplicate the dataset and perform cross-scale sampling based on error rates. This facilitates cross-error joint training with a mix of problem difficulties.
  3. 3.
    DPO Training on the Aligned Model: Building on the format-aligned model, we further train using a DPO approach.
This phased approach achieves significant capability recovery while introducing less than 1% additional parameter updates relative to the original model sizes. Our experiments demonstrate that targeted restoration using as few as 545 error-specific examples per category can bridge the performance gap between quantized and full-precision models. Remarkably, even training with a dataset comprising only 323 samples for a single error type can restore up to 9% of the model’s capability, confirming that localized capability losses can be efficiently remediated through structured retraining.
Algorithm 1 Pseudocode of our approach.
Input: Full-precision model $M$, Quantization method $Q$, Step-annotated dataset $D_{train}$, Evaluation dataset $D_{test}$, Expert models $\{E_{1}...E_{n}\}$
Output: Enhanced quantized model $M_{enhanced}$
1:  Step 1: Model Quantization
2:  Apply quantization method $Q$ to model $M$:
3:  $M_{Q}\leftarrow Q(M)$
4:  Step 2: Format Alignment
5:  Fine-tune $M_{Q}$ with QLoRA on $D_{train}$:
6:  $M_{A}\leftarrow\text{QLoRA}(M_{Q},D_{train})$
7:  Step 3: Error Evaluation
8:  Generate outputs $O$ on $D_{test}$:
9:  $O\leftarrow M_{A}(D_{test})$
10:  Initialize error set $E\leftarrow\emptyset$
11:  for each output $o_{i}\in O$ do
12:     Multi-Expert Evaluation:
13:     Collect evaluations $\{E_{1}(o_{i})...E_{n}(o_{i})\}$
14:     if Majority vote $\neq$ DeepSeek-R1 and votes $<$then
15:        Flag $o_{i}$ as Rejected
16:        $E\leftarrow E\cup\{\text{ErrorType}(o_{i})\}$
17:     end if
18:  end for
19:  Step 4: Human Verification
20:  Random sample $s\leftarrow 2\%$ of $E$:
21:  $E_{sample}\leftarrow\text{RandomSample}(E,s)$
22:  $E_{verified}\leftarrow\text{HumanVerify}(E_{sample})$
23:  Step 5: Capability Restoration
24:  for each error type $t\in E_{verified}$ do
25:     Generate targeted dataset $D_{t}$:
26:     $D_{t}\leftarrow\text{CurateData}(t,D_{train})$
27:     Apply DPO training:
28:     $M_{A}\leftarrow\text{DPO}(M_{A},D_{t})$
29:  end for
30:  Return: $M_{enhanced}\leftarrow M_{A}$
4 Experiments
4.1 General Setup
All experiments were conducted on a cluster comprising 4 Nvidia 80G A800 GPUs.
Base Model & Scale
Given the general applicability of our approach to mainstream models and the versatility across different model sizes, we select three distinct models from the Llama 3 series as the foundation for our experiments. Specifically, we choose models of varying sizes, namely 8B for Llama-3.1 and 1B/3B for Llama-3.2, all Instruct version.
For all of these models, we observe significant math accuracy degradation resulting from quantization, and we also demonstrate the powerful effect of our proposed fast reconstruction method.
Quantization Strategy. Our experiments employ two mainstream quantization techniques: AWQ and GPTQ Frantar et al. (2022); Lin et al. (2024), both known for their strong performance in model compression. These methods have led to numerous low-precision models widely adopted in practice. Our work aims to provide complementary insights for community.
Training Strategy. We use a sequence length of 1024 and in both the format alignment and capability restoration stages, we employed a global batch size of 16 and a warm-up period spanning 3%. All models were trained for 3 epochs.
Evaluation Strategy. We evaluated our models using the OpenCompass evaluation framework Contributors (2023), maintaining consistency in both the global batch size and the prompt with those used during training. This setup ensures a fair comparison across all models.
4.2 Format Alignment
Setting. We applied LoRA fine-tuning to the vanilla model, while QLoRA Dettmers et al. (2024) was employed for the quantized models. The LoRA rank was set to 32 for the format alignment stage. During training, the system prompt strictly enforced a step-wise output of the reasoning process, and outputs that deviated from the required format were penalized.
Figure 5: Comparative analysis of error types for quantized models (AWQ-W4A16 and GPTQ-W4A16) across three model scales.
4.3 Analyses for Reasoning Errors
Error Assessment Pipeline. To facilitate a rigorous and scalable evaluation of quantization-induced errors in reasoning tasks, we developed an automated assessment pipeline powered by state-of-the-art language models. This pipeline systematically processes model outputs and classifies errors according to our predefined error_types_list taxonomy. By leveraging a pre-trained transformer as the core evaluator, we reduce subjective bias and ensure consistent, reproducible error analyses across all experimental conditions. Furthermore, the computational scoring framework supports high-throughput performance assessment while preserving granularity in error categorization.
Our evaluation pipeline comprises three primary stages:
  1. 1.
    Model Judge: For each instance in which a quantized model produces an incorrect answer, we employ a dedicated “judge model” to analyze the error. The judge model is tasked with: (a) identifying the first occurrence of an error, (b) specifying the exact step where the error is introduced, (c) assigning an error category based on a nested classification scheme, and (d) providing an explanation along with a confidence score for its determination.
  2. 2.
    Majority Vote: To mitigate hallucinations and enhance evaluation stability, we implement a majority voting mechanism across five models: DeepSeek-R1, GPT-4o, GPT-4, Qwen-Max, and DeepSeek-V3. DeepSeek-R1’s assessment is treated as the baseline, and its results are compared with the majority vote from the other models. Instances of disagreement are flagged for further review, ensuring consistency and minimizing spurious judgments.
  3. 3.
    Human Verification: For cases with conflicting assessments from the majority vote, manual review is conducted. Additionally, we randomly sample 2% of the evaluated cases to verify the accuracy and consistency of the automated judgments.
[…]
MATH-500
@@ 4.3 Analyses for Reasoning Errors
Llama-3.1-8B-Instruct Aligned score 37.36 30.4 30.8
$\Delta$downarrow6.94(18.58%) $\downarrow$6.56(17.56%)
Llama-3.2-3B-Instruct Aligned score 37.4 34.4 31.4
$\Delta$downarrow3.0(8.02%) $\downarrow$6.0(16.04%)
Llama-3.2-1B-Instruct Aligned score 21 14.2 14.94
$\Delta$downarrow6.8(32.39%) $\downarrow$6.2(29.52%)
Table 2: Evaluation results of format-aligned full-precision models and quantization methods (AWQ and GPTQ) on the MATH-500 benchmark. After alignment, the model is required to output responses strictly following the numbered step-wise format, resulting in slight changes in evaluation score.
@@
Quantitative analyses. Table 1 presents a performance comparison between full-precision and quantized models on the MATH dataset across three model scales. We report both the performance (Score) and the corresponding accuracy degradation ($\Delta$) induced by quantization. The vanilla model serves as the upper baseline, while quantized models exhibit varying degrees of performance loss. Notably, our results indicate that the 1B model is particularly sensitive to quantization.
Based on the set of questions answered incorrectly by the quantized models, we conducted further quantitative analyses. Figure 3 and 5 illustrates the overall distribution of the four error categories. Conceptual errors account for the majority, representing more than half of all observed mistakes. Execution errors follow next. Our case study analyses reveal that quantized models are particularly deficient in basic arithmetic operations, with multiplication errors showing the most significant performance decline. We hypothesize that these computation errors, especially in models quantized to 4-bit weights, are likely attributable to low-bit precision issues such as overflow and underflow, which propagate inaccuracies throughout multi-step calculations. Additional illustrative cases are provided in the Appendix.
@@
Intuitively, one might expect that reasoning errors would be more frequent given the inherent complexity of mathematical reasoning. However, our analysis shows that reasoning errors are less common. This observation may reflect a “survivor effect,” whereby earlier errors, such as logical or computational failures, prevent the model from reaching subsequent steps where boundary conditions would be tested. For precise statistics, we recorded only the first error type encountered in each instance. These findings suggest that quantization disrupts the step-by-step reasoning process, thereby exposing the models’ limitations in managing complex mathematical tasks. The underlying causes likely stem from limited numerical precision and the accumulation of quantization noise in low-bit formats, which impair the representation of long-term dependencies and intermediate values during complex reasoning.
We hypothesize that small losses in specific capability dimensions can be compensated by fine-tuning on a minimal subset of targeted data. This approach could potentially restore overall performance by addressing the specific weaknesses introduced by quantization. For instance, targeted training on multi-step reasoning tasks could reduce step omissions by reinforcing the model’s ability to retain intermediate results. By strategically addressing these specific failure modes, it may be possible to recover and enhance the overall capabilities of quantized models, leveraging their efficiency while minimizing performance degradation.
Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct
Full Precision 18.6 40.2
AWQ GPTQ AWQ GPTQ
Aligned 14.26 14.94 38.8 31.6
Setting 0 17.2 18.6 41.6 38.2
Setting 1 15.8 18.0 41.2 36.2
Setting 2 16.6 16.6 37.2 34.4
Setting 3 15.2 15.4 35.2 33.0
Table 3: Ablation study settings and evaluation results on the MATH-500 dataset. Setting 0: All error cases; Setting 1: Conceptual error cases only; Setting 2: Methodological error cases only; Setting 3: Execution error cases only.
4.4 Capability Restore
Dataset Sampling. We constructed the dataset for capability restoration by sampling from all MATH problems that were answered correctly by the full-precision model but incorrectly by the quantized model. Data sampling was guided by proportional representation and consistency with expert model major vote outcomes. To enhance data validity, we performed deduplication and prioritized cases where the major vote results were most consistent and the errors were more pronounced. We aimed to maintain a compact dataset to reduce training resource consumption. Ultimately, we effectively sampled 545 examples from 3329 quantized model failure cases. The full-precision model’s correct answers were used as positive examples, while the quantized model’s incorrect answers served as negative examples, forming our DPO dataset.
Setting. Training for capability restoration stage was conducted on the format-aligned model. For this stage, the LoRA rank was still set to 32, and a cosine learning rate schedule was employed with a warm-up ratio of 0.1. We used a sigmoid function as the preference loss function, with a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For evaluation, to ensure fairness and mitigate data leakage risks, we used the MATH-500 dataset. MATH-500 comprises 500 problems that are not included in the PRM800K dataset and were randomly sampled from the MATH dataset, thereby ensuring a similar difficulty distribution.
Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct
Full Precision 21 37.4
AWQ GPTQ AWQ GPTQ
Format Aligned 14.26 14.94 38.8 31.6
Restored 18.2 18.8 41.6 38.2
Table 4: Capability restoration experiment results on the MATH-500 benchmark. Using only 545 samples within a few minutes.
Results. We conducted rapid capability restoration training on two smaller-scale models. Table 4 presents the restoration results using a dataset of 545 sampled examples. Due to the limited data volume, training required only 2–4 minutes on 4 GPUs, demonstrating minimal resource consumption. Notably, the quantized models exhibited marked improvements on this dataset.
4.5 Ablation Study
We conducted quantitative analyses focusing on the earliest error type observed during each model’s reasoning process. We hypothesize that quantized models may exhibit a “bucket effect” in reasoning tasks, wherein they still perform well on instruction following yet fail to solve complex reasoning problems. To investigate this, we isolated the three most frequently occurring early error types for ablation experiments: (1) Conceptual error cases only, (2) Method error cases only, (3) Execution errors only. (4) All error types cases. This targeted analysis allowed us to assess the specific contributions of each error category to the overall degradation in reasoning performance, thereby providing insights into how quantization selectively impacts different aspects of model capability.
The ablation experiments results reveal that even when training on datasets containing only a single type of error, the model’s capabilities can be significantly improved. For instance, using only data associated with conceptual errors enabled the quantized model to recover reasoning performance comparable to that of the full-precision model. Notably, this subset comprised merely 323 examples and takes very few minutes for training, which required minimal computational resources. These findings lend support to our hypothesis: quantization does not diminish the model’s fundamental abilities but selectively degrades specific dimensions that are crucial for solving reasoning tasks, such as those encountered in mathematical problem solving.